OpenStudy (anonymous):

now i have 20*10*5 thick plate ... i want 5000 pieces how can calculating in plate

OpenStudy (anonymous):

not getting the question:((((

OpenStudy (shane_b):

I don't really understand the question either but maybe this will help. I'll assume the lengths are measured in meters: Calculate the total volume of the plate:\[V=LWH=20m*10m*5m=1000m^3\]You want to cut that into 5000 pieces so just divide the total volume by 5000:\[\frac{1000m^3}{5000}=0.2m^3\]So each piece would need to be 0.2m^3 in volume. To get the L, W, and H of each cube simply take the cube root of 0.2m^3.\[\sqrt[3]{0.2m^3}=0.585m\]So the length, width and height of each piece would be 0.585m.